# On model emulation and closure tests for 3+1D relativistic heavy-ion collisions

### v2

@@ I Introduction
However, direct measurement of the resultant hot and dense QGP state remains elusive due to its transient existence and confinement to small spatial dimensions within collider experiments. Understanding the QGP’s properties necessitates analyzing experimental measurements, which imprint the entire collision history as the QGP is created, expands, and fragments into a collection of hadrons Busza et al. (2018). Theoretical frameworks for modeling such collisions have undergone substantial advancement over recent decades, encompassing various stages from the initial state Miller et al. (2007); Loizides et al. (2015); Broniowski et al. (2009); Drescher and Nara (2007); Schenke et al. (2012); Paatelainen et al. (2013); Moreland et al. (2015); Giacalone et al. (2019); Shen and Schenke (2018) to hadronic rescattering dynamics Bass et al. (1998); Weil et al. (2016); Cassing and Bratkovskaya (1999); Lin et al. (2005), including pre-equilibrium Schlichting and Teaney (2019), relativistic viscous hydrodynamics Teaney (2010); Romatschke and Romatschke (2019), and particlization Cooper and Frye (1974).
@@
Bayesian inference, leveraging Bayes’s theorem, offers a principled approach to address inverse problems by discerning parameter sets and their associated uncertainties systematically Sivia and Skilling (2006). However, the exhaustive evaluation of models across diverse parameter spaces often proves infeasible due to computational constraints. Herein lies the utility of model emulation, which, once trained with model outputs, furnish computationally inexpensive surrogates for the full model. These emulators, by virtue of their efficiency, replace the full model in Bayesian analyses, thereby facilitating parameter calibration.
@@ II Bayesian Inference
Multi-stage event-by-event simulation frameworks, such as iEBE-VISHNU Shen et al. (2016); Bernhard et al. (2016), JETSCAPE Putschke et al. (2019), Trajectum Nijs et al. (2021a), and iEBE-MUSIC Shen and Schenke (2022), are essential tools for studying various physics phenomena in relativistic heavy-ion collisions. However, these simulations are often computationally intensive. To alleviate this computational burden, emulators are employed to predict the model output for a given parameter set Kennedy and O’Hagan (2001); Higdon et al. (2004).
@@
The problem at hand is inherently an inverse problem, where we seek to identify the parameters $\bm{\theta}$ that best describe the observed data. To address this, we adopt a Bayesian approach to inference, where probability is interpreted as the degree of belief accounting for all available information Sivia and Skilling (2006). Bayes’ rule provides a framework for updating our beliefs based on observed data:
@@
To obtain the posterior distribution, which is often intractable analytically, Markov Chain Monte Carlo (MCMC) techniques are employed Brooks et al. (2011); Trotta (2008). These methods enable the estimation of posterior distributions for $\bm{\theta}$ and quantifying their uncertainties.
In practice, evaluating the right-hand side of Eq. (2) for numerous $\bm{\theta}$ sets in MCMC simulations can be computationally prohibitive. Here, GP emulators offer a solution by providing computationally efficient surrogates for the expensive simulations Gramacy (2020); Rasmussen (2004). These emulators are trained using simulation data for various parameter sets $\bm{\theta}$. Once trained, the emulators can predict the mean $\bm{\mu}(\bm{\theta})$ and covariance ${\bf C}(\bm{\theta})$ of the simulation output ${\bf y}_{\rm exp}(\bm{\theta})$. Subsequently, the likelihood function (Eq. (3)) can be approximated as Kennedy and O’Hagan (2001):
@@ III Gaussian Process Emulators
This section presents a comparative analysis of different implementations of GP emulators employed in our study. Specifically, we investigate two GP emulator methodologies, namely the PCGP and the PCSK methods, which are integrated into the Python package surmise by the BAND collaboration Plumlee et al. (2023). Additionally, we consider the standard GP emulator available in the Scikit-learn Python package. The Scikit GP was used in a few previous Bayesian analyses in our field Bernhard et al. (2019); Everett et al. (2021a); Mäntysaari et al. (2022); Shen et al. (2024, 2023).
All these emulators share a basis vector approach Higdon et al. (2008), tailored to address the challenges posed by the high-dimensional output space of the model. Principal Component Analysis (PCA) Ramsay and Silverman (1997) is utilized to project the high-dimensional output onto a lower-dimensional space, enabling efficient emulation. Specifically, each output for a principal component is independently trained by a GP emulator.
We begin by elucidating the fundamental concept behind GP emulation with PCA. Let $\{\bm{\theta}_{1}^{\text{tr}},\dots,\bm{\theta}_{n}^{\text{tr}}\}$ represent a training parameter set comprising $n$ unique training points for the emulators. Each training point $\bm{\theta}_{i}^{\text{tr}}$ generates a $d$-dimensional vector $\bar{\mathbf{y}}_{\text{sim}}(\bm{\theta}_{i}^{\text{tr}})$, where $d$ denotes the number of observables. The model simulation results $\bar{\mathbf{y}}_{\text{sim}}(\bm{\theta}_{i}^{\text{tr}})$ are averaged over multiple collision events. Before applying the PCA transformation, we standardize the training data as, $\tilde{\bf y}_{\rm sim}(\bm{\theta}_{i}^{\rm tr})\equiv\mathbf{G}^{-1}(\bar{\bf y}_{\rm sim}(\bm{\theta}_{i}^{\rm tr})-\bf h)$, using a centering vector $\mathbf{h}$ and scaling matrix $\mathbf{G}=\mathrm{diag}(\bf c)$. The resulting $\tilde{\bf y}_{\rm sim}(\bm{\theta}_{i}^{\rm tr})$ has a zero mean and unity variance over the $n$ training samples. We define a $d\times n$ matrix $\mathbf{\Xi}\equiv\{\tilde{\mathbf{y}}_{\text{sim}}(\bm{\theta}_{1}^{\text{tr}}),\dots,\tilde{\mathbf{y}}_{\text{sim}}(\bm{\theta}_{n}^{\text{tr}})\}$. PCA yields a linear transformation matrix ${\bf S}=\left[{\bf s}_{1},\dots,{\bf s}_{p}\right]$ of dimension $d\times p$, which can project the output matrix $\bm{\Xi}$ from a $d$-dimensional space to a $p$-dimensional principal component space via the transformation ${\bf T}=\left[{\bf t}_{1},\dots,{\bf t}_{p}\right]=\bf S^{\mathsf{T}}\bm{\Xi}$. In most applications, keeping only $p\ll d$ principal components to describe almost all the variance in the original dataset is sufficient Higdon et al. (2015); Sürer et al. (2022); Liyanage et al. (2023); Moses Y.-H. Chan and Wild (2024).
@@
where $m_{l}(\bm{\theta})=\mathbf{k}_{l}^{\mathsf{T}}\mathbf{K}_{l}^{-1}\mathbf{t}_{l}$ and $s^{2}_{l}(\bm{\theta})=k_{l}(\bm{\theta},\bm{\theta})-\mathbf{k}_{l}^{\mathsf{T}}(\bm{\theta})\mathbf{K}_{l}^{-1}\mathbf{k}_{l}(\bm{\theta})$. Here, $\mathbf{k}_{l}(\bm{\theta})=\left[k_{l}(\bm{\theta},\bm{\theta}^{\rm tr}_{i})\right]^{n}_{i=1}$ denotes the covariance vector between $n$ training parameters $\{\bm{\theta}_{1}^{\rm tr},\dots,\bm{\theta}_{n}^{\rm tr}\}$ and any chosen parameter $\bm{\theta}$, and $\mathbf{K}_{l}=\left[k_{l}(\bm{\theta}_{i}^{\rm tr},\bm{\theta}_{j}^{\rm tr})+\delta_{i,j}r_{l,i}\right]_{i,j=1}^{n}$ represents the covariance matrix between the $n$ training parameters. The covariance function $k_{l}(\bm{\theta},\bm{\theta}^{\prime})$ depends on the choice of the kernel function in the GP. [^1] The PCSK and PCGP emulators use a Matérn kernel (smoothness parameter $3/2$), while we choose to use a Radial Basis Function (RBF) kernel for the Scikit GP. The Scikit GP setup was used in a few of the previous Bayesian analyses Everett et al. (2021a); Shen et al. (2024). Additionally, the term $\delta_{i,j}r_{l,i}$, along with the Kronecker delta, accounts for uncertainty arising from stochastic simulation results for the different events. Here $r_{l,i}$ is the square of the statistical uncertainty of the $l$-th principal component at the training point $\bm{\theta}_{i}$.
The term $\delta_{i,j}r_{l,i}$ is the main difference between the PCGP/Scikit GP implementations of GPs used in almost all heavy-ion physics related studies Petersen et al. (2011); Novak et al. (2014); Sangaline and Pratt (2016); Bernhard et al. (2015, 2016); Moreland et al. (2020); Bernhard et al. (2019); Everett et al. (2021b); Nijs et al. (2021b, a); Everett et al. (2021a); Heffernan et al. (2023a, b) and the PCSK method used in Liyanage et al. (2023). Instead of only taking into account the mean values of the training data, the PCSK method also considers the standard deviation at the training points. This is particularly relevant as simulations of heavy-ion collisions with different initial states for given parameters $\bm{\theta}$ yield observables with a certain variation. The PCSK emulation incorporates this variation in the accuracy between the resulting data points, known as heteroskedasticity, by using stochastic kriging Ankenman et al. (2010). We explore the effects of the size of statistical errors on the trained emulator accuracy in Appendix A.
@@ IV.1 Training Setup
In this work, we divide the experimental observables into several groups and apply the principal component Gaussian Process Regression for each group Liyanage et al. (2023); Shen et al. (2024). In our work, we train GP emulators for different observable groups first separated by the system’s collision energy. We further split the particle multiplicities observables from the anisotropic flow coefficients and average transverse momenta because we will test the GP emulators’ accuracy for training with multiplicities or the logarithm of the data. The logarithmic transformation was adopted by a few previous works Everett et al. (2021a); Nijs et al. (2021a).
@@
The datasets contain event-by-event simulations of Au+Au collisions at three energies: 200, 19.6, and 7.7 GeV. The training dataset contains 1000 design points determined with a Maximum Projection Latin Hypercube Design algorithm Roshan Joseph et al. (2020, 2015) to explore the $\bm{\theta}$-space. This algorithm orders the design points to ensure uniform coverage of model parameter space with a subset of design points. This feature allows us to quantify how the emulator accuracy scales with the size of training datasets. In the following, we denote the design points from this algorithm by “LHD points”. For each design point, 1000 minimum bias events were simulated at 200 GeV and 2000 minimum bias events at 19.6 and 7.7 GeV each. [^2] Points with relative statistical uncertainty larger than 10% are discarded before training the GP emulators. Additionally, 100 parameter sets were drawn from the resulting posterior distributions in Ref. Shen et al. (2024), and simulations were carried out for these points. This allows us to study, in an active learning scenario, whether the uncertainty in the emulators can be decreased when the density of training points is increased in regions where the likelihood of the final parameters is more significant.
@@ IV.2 PCA Transformation for Functional Model Parameters
[…]
@@
[…]
@@
Although the model parameters in Table 2 provide plenty of information to the GP emulators, the functional form of the parametrizations is not included in the training. One way to provide the GP emulators with this information is to enlarge the dimension of the model parameter space and feed the GP emulators with more sampled points for $\langle y_{\rm loss}\rangle(y_{\mathrm{init}})$ and QGP specific shear and bulk viscosities $\tilde{\eta}(\mu_{B})$ and $\tilde{\zeta}(T,\mu_{B})$. This approach was carried out in the previous JETSCAPE analysis and effectively increased emulator accuracy Everett et al. (2021a, b). Here, we introduce a more systematic way to deal with functional model parameters using PCA transformation.
@@ IV.4.2 Benchmark Emulation Performance
Figure 2 displays the emulators’ RMS error $\mathcal{E}$ for all 544 experimental observables – separated by black lines according to Table 1 – trained with the three types of GP emulators. The upper panel summarizes the mid-rapidity observables from central to 60% peripheral centrality for Au+Au collisions at three energies. The lower panel shows the pseudo-rapidity dependence observables in five centrality classes at 19.6 and 200 GeV. For the ${\rm d}N_{\rm ch}/{\rm d}\eta$ datasets, we can identify the different “$w$”-shaped structures, which are the multiplicities as a function of rapidity in different centrality classes. Each point in the plot represents a single observable, indicated by the ticks on the $x$-axis.
@@
Overall, we find the PCGP and PCSK emulators from the BAND surmise package perform better than the Scikit GP emulator with standard settings. We note that the relatively poorer performance of the Scikit GP emulator could be related to not having optimized the GP kernel and its parameters.

### v3

@@ I Introduction
However, direct measurement of the resultant hot and dense QGP state remains elusive due to its transient existence and confinement to small spatial dimensions within collider experiments. Understanding the QGP’s properties necessitates analyzing experimental measurements, which imprint the entire collision history as the QGP is created, expands, and fragments into a collection of hadrons Busza et al. (2018). Theoretical frameworks for modeling such collisions have undergone substantial advancement over recent decades, encompassing various stages from the initial state Miller et al. (2007); Loizides et al. (2015); Broniowski et al. (2009); Drescher and Nara (2007); Schenke et al. (2012); Paatelainen et al. (2013); Moreland et al. (2015); Giacalone et al. (2019); Shen and Schenke (2018) to hadronic rescattering dynamics Bass and others (1998); Weil and others (2016); Cassing and Bratkovskaya (1999); Lin et al. (2005), including pre-equilibrium Schlichting and Teaney (2019), relativistic viscous hydrodynamics Teaney (2010); Romatschke and Romatschke (2019), and particlization Cooper and Frye (1974).
@@
Bayesian inference, leveraging Bayes’s theorem, offers a principled approach to address inverse problems by discerning parameter sets and their associated uncertainties systematically Sivia2006. However, the exhaustive evaluation of models across diverse parameter spaces often proves infeasible due to computational constraints. Herein lies the utility of model emulation, which, once trained with model outputs, furnishes computationally inexpensive surrogates for the full model. These emulators, by virtue of their efficiency, replace the full model in Bayesian analyses, thereby facilitating parameter calibration.
@@ II Bayesian Inference
Multi-stage event-by-event simulation frameworks, such as iEBE-VISHNU Shen et al. (2016); Bernhard et al. (2016), JETSCAPE Putschke and others (2019), Trajectum Nijs et al. (2021a), and iEBE-MUSIC Shen and Schenke (2022), are essential tools for studying various physics phenomena in relativistic heavy-ion collisions. However, these simulations are often computationally intensive. To alleviate this computational burden, emulators are employed to predict the model output for a given parameter set Ohagan2001; Higdon2004.
@@
The problem at hand is inherently an inverse problem, where we seek to identify the parameters $\bm{\theta}$ that best describe the observed data. To address this, we adopt a Bayesian approach to inference, where probability is interpreted as the degree of belief accounting for all available information Sivia2006. Bayes’ rule provides a framework for updating our beliefs based on observed data:
@@
To obtain the posterior distribution, which is often intractable analytically, Markov Chain Monte Carlo (MCMC) techniques are employed brooks2011handbook; Trotta2008. These methods enable the estimation of posterior distributions for $\bm{\theta}$ and quantifying their uncertainties.
In practice, evaluating the right-hand side of Eq. (2) for numerous $\bm{\theta}$ sets in MCMC simulations can be computationally prohibitive. Here, GP emulators offer a solution by providing computationally efficient surrogates for the expensive simulations gramacy2020surrogates; Rasmussen2004. These emulators are trained using simulation data for various parameter sets $\bm{\theta}$. Once trained, the emulators can predict the mean $\bm{\mu}(\bm{\theta})$ and covariance ${\bf C}(\bm{\theta})$ of the simulation output ${\bf y}_{\rm exp}(\bm{\theta})$. Subsequently, the likelihood function (Eq. (3)) can be approximated as Ohagan2001:
@@ III Gaussian Process Emulators
This section presents a comparative analysis of different implementations of GP emulators employed in our study. Specifically, we investigate two GP emulator methodologies, namely the PCGP and the PCSK methods, which are integrated into the Python package surmise by the BAND collaboration surmise2023. Additionally, we consider the standard GP emulator available in the Scikit-learn Python package. The Scikit GP was used in a few previous Bayesian analyses in our field Bernhard et al. (2019); Everett and others (2021a); Nijs et al. (2021b); Parkkila et al. (2021); Mäntysaari et al. (2022); Heffernan et al. (2023b); Soeder et al. (2023); Shen et al. (2024, 2023).
All these emulators share a basis vector approach Higdon2008, tailored to address the challenges posed by the high-dimensional output space of the model. Principal Component Analysis (PCA) Ramsay97functionaldata is utilized to project the high-dimensional output onto a lower-dimensional space, enabling efficient emulation. Specifically, each output for a principal component is independently trained by a GP emulator.
We begin by elucidating the fundamental concept behind GP emulation with PCA. Let $\{\bm{\theta}_{1}^{\text{tr}},\dots,\bm{\theta}_{n}^{\text{tr}}\}$ represent a training parameter set comprising $n$ unique training points for the emulators. Each training point $\bm{\theta}_{i}^{\text{tr}}$ generates a $d$-dimensional vector $\bar{\mathbf{y}}_{\text{sim}}(\bm{\theta}_{i}^{\text{tr}})$, where $d$ denotes the number of observables. The model simulation results $\bar{\mathbf{y}}_{\text{sim}}(\bm{\theta}_{i}^{\text{tr}})$ are averaged over multiple collision events. Before applying the PCA transformation, we standardize the training data as, $\tilde{\bf y}_{\rm sim}(\bm{\theta}_{i}^{\rm tr})\equiv\mathbf{G}^{-1}(\bar{\bf y}_{\rm sim}(\bm{\theta}_{i}^{\rm tr})-\bf h)$, using a centering vector $\mathbf{h}$ and scaling matrix $\mathbf{G}=\mathrm{diag}(\bf c)$. The resulting $\tilde{\bf y}_{\rm sim}(\bm{\theta}_{i}^{\rm tr})$ has a zero mean and unity variance over the $n$ training samples. We define a $d\times n$ matrix $\mathbf{\Xi}\equiv\{\tilde{\mathbf{y}}_{\text{sim}}(\bm{\theta}_{1}^{\text{tr}}),\dots,\tilde{\mathbf{y}}_{\text{sim}}(\bm{\theta}_{n}^{\text{tr}})\}$. PCA yields a linear transformation matrix ${\bf S}=\left[{\bf s}_{1},\dots,{\bf s}_{p}\right]$ of dimension $d\times p$, which can project the output matrix $\bm{\Xi}$ from a $d$-dimensional space to a $p$-dimensional principal component space via the transformation ${\bf T}=\left[{\bf t}_{1},\dots,{\bf t}_{p}\right]=\bf S^{\mathsf{T}}\bm{\Xi}$. In most applications, keeping only $p\ll d$ principal components to describe almost all the variance in the original dataset is sufficient Higdon et al. (2015); Sürer et al. (2022); Liyanage et al. (2023); doi:10.1080/00401706.2023.2210170.
@@
where $m_{l}(\bm{\theta})=\mathbf{k}_{l}^{\mathsf{T}}\mathbf{K}_{l}^{-1}\mathbf{t}_{l}$ and $s^{2}_{l}(\bm{\theta})=k_{l}(\bm{\theta},\bm{\theta})-\mathbf{k}_{l}^{\mathsf{T}}(\bm{\theta})\mathbf{K}_{l}^{-1}\mathbf{k}_{l}(\bm{\theta})$. Here, $\mathbf{k}_{l}(\bm{\theta})=\left[k_{l}(\bm{\theta},\bm{\theta}^{\rm tr}_{i})\right]^{n}_{i=1}$ denotes the covariance vector between $n$ training points $\{\bm{\theta}_{1}^{\rm tr},\dots,\bm{\theta}_{n}^{\rm tr}\}$ and any chosen parameter vector $\bm{\theta}$, and $\mathbf{K}_{l}=\left[k_{l}(\bm{\theta}_{i}^{\rm tr},\bm{\theta}_{j}^{\rm tr})+\delta_{i,j}r_{l,i}\right]_{i,j=1}^{n}$ represents the covariance matrix between the $n$ training points. The covariance function $k_{l}(\bm{\theta},\bm{\theta}^{\prime})$ depends on the choice of the kernel function in the GP. [^1] The PCSK and PCGP emulators use a Matérn kernel (smoothness parameter $3/2$), while we choose to use a Radial Basis Function (RBF) kernel for the Scikit GP. The Scikit GP setup was used in a few of the previous Bayesian analyses Everett and others (2021a); Shen et al. (2024). Additionally, the term $\delta_{i,j}r_{l,i}$, along with the Kronecker delta, accounts for uncertainty arising from stochastic simulation results for the different events. Here $r_{l,i}$ is the square of the statistical uncertainty of the $l$-th principal component at the training point $\bm{\theta}_{i}$.
The term $\delta_{i,j}r_{l,i}$ is the main difference between the PCGP/Scikit GP implementations of GPs used in almost all heavy-ion physics related studies Petersen et al. (2011); Novak et al. (2014); Sangaline and Pratt (2016); Bernhard et al. (2015, 2016); Moreland et al. (2020); Bernhard et al. (2019); Everett and others (2021b); Nijs et al. (2021b, a); Everett and others (2021a); Heffernan et al. (2023a, b) and the PCSK method used in Liyanage et al. (2023). Instead of only taking into account the mean values of the training data, the PCSK method also considers the standard deviation at the training points. This is particularly relevant as simulations of heavy-ion collisions with different initial states for given parameters $\bm{\theta}$ yield observables with a certain variation. The PCSK emulation incorporates this variation in the accuracy between the resulting data points, known as heteroskedasticity, by using stochastic kriging Ankenman2010. We explore the effects of the size of statistical errors on the trained emulator accuracy in Appendix A.
@@ IV.1 Training Setup
In this work, we divide the experimental observables into several groups and apply the principal component Gaussian Process Regression for each group Liyanage et al. (2023); Shen et al. (2024). In our work, we train GP emulators for different observable groups first separated by the system’s collision energy. We further split the particle multiplicities observables from the anisotropic flow coefficients and average transverse momenta because we will test the GP emulators’ accuracy for training with multiplicities or the logarithm of the data. The logarithmic transformation was adopted by a few previous works Everett and others (2021a); Nijs et al. (2021a).
@@
The datasets contain event-by-event simulations of Au+Au collisions at three energies: 200, 19.6, and 7.7 GeV. The training dataset contains 1000 design points determined with a Maximum Projection Latin Hypercube Design algorithm doi:10.1080/00224065.2019.1611351; 10.1093/biomet/asv002 to explore the $\bm{\theta}$-space. This algorithm orders the design points to ensure uniform coverage of model parameter space with a subset of design points. This feature allows us to quantify how the emulator accuracy scales with the size of training datasets. In the following, we denote the design points from this algorithm by “LHD points”. For each design point, 1000 minimum bias events were simulated at 200 GeV and 2000 minimum bias events at 19.6 and 7.7 GeV each. [^2] Points with relative statistical uncertainty larger than 10% are discarded before training the GP emulators. Additionally, 100 parameter sets were drawn from the resulting posterior distributions in Ref. Shen et al. (2024), and simulations were carried out for these points. This allows us to study, in an active learning scenario, whether the uncertainty in the emulators can be decreased when the density of training points is increased in regions where the likelihood of the final parameters is more significant.
@@ IV.2 PCA Transformation for Functional Model Parameters
$\langle y_{\rm loss}\rangle=\begin{cases}y_{\rm loss,2}\frac{y_{\rm init}}{2},\;0<y_{\rm init}\leq 2\\ y_{\rm loss,2}+(y_{\rm loss,4}-y_{\rm loss,2})\frac{y_{\rm init}-2}{2},\;2<y_{\rm init}<4\\ y_{\rm loss,4}+(y_{\rm loss,6}-y_{\rm loss,4})\frac{y_{\rm init}-4}{2},\;y_{\rm init}\geq 4\end{cases}$ (7)
@@
$\tilde{\eta}(\mu_{B})=\begin{cases}\eta_{0}+(\eta_{2}-\eta_{0})\frac{\mu_{B}}{0.2},\;0<\mu_{B}\leq 0.2\;\mathrm{GeV}\\ \eta_{2}+(\eta_{4}-\eta_{2})\frac{(\mu_{B}-0.2)}{0.2},\;0.2<\mu_{B}<0.4\;\mathrm{GeV}\\ \eta_{4},\;\mu_{B}\geq 0.4\;\mathrm{GeV}\end{cases}$ (8)
@@
Although the model parameters in Table 2 provide plenty of information to the GP emulators, the functional form of the parametrizations is not included in the training. One way to provide the GP emulators with this information is to enlarge the dimension of the model parameter space and feed the GP emulators with more sampled points for $\langle y_{\rm loss}\rangle(y_{\mathrm{init}})$ and QGP specific shear and bulk viscosities $\tilde{\eta}(\mu_{B})$ and $\tilde{\zeta}(T,\mu_{B})$. This approach was carried out in the previous JETSCAPE analysis and effectively increased emulator accuracy Everett and others (2021a, b). Here, we introduce a more systematic way to deal with functional model parameters using PCA transformation.
@@ IV.4.2 Benchmark Emulation Performance
Figure 2 displays the emulators’ RMS error $\mathcal{E}$ for all 544 experimental observables – separated by black lines according to Table 1 – trained with the three types of GP emulators. The upper panel summarizes the mid-rapidity observables from central to 60% peripheral centrality for Au+Au collisions at three energies. The lower panel shows the pseudo-rapidity dependence observables in five centrality classes at 19.6 and 200 GeV. For the ${\rm d}N_{\rm ch}/{\rm d}\eta$ datasets, we can identify the different “ww”-shaped structures, which are the multiplicities as a function of rapidity in different centrality classes. Each point in the plot represents a single observable, indicated by the ticks on the $x$-axis.
@@
Overall, we find the PCGP and PCSK emulators from the BAND surmise package perform better than the Scikit GP emulator with standard settings. We note that the relatively poorer performance of the Scikit GP emulator could be related to not having optimized the GP kernel and its parameters. [^4] The Scikit GP kernel consists of an RBF kernel with an additional white noise term similar to Eq. (4.29) in Bernhard (2018). The choice of default kernel parameters followed the JETSCAPE calibration Everett and others (2021a).